# Janus-faces of Temporal Constraint Languages: A dichotomy of expressivity

### v1

@@ 2.2. Graph-theoretic
A digraph is a relational structure of the form $\mathbb{G}=(G;E)$, where $E$ is binary. $\mathbb{G}$ is called smooth if $\textnormal{pr}_{1}(E)=\textnormal{pr}_{2}(E)$. If $E$ is symmetric, $\mathbb{G}$ is also called a graph. For $m\in\mathbb{N}$, we denote by $E^{m}$ the binary relation containing all tuples $(a_{0},a_{m})\in G\times G$ for which there exist $a_{1},\dots,a_{m-1}$ such that $(a_{i-1},a_{i})\in E$ for all $i\leq m$. We write $E^{-1}$ for the relation $\{(y,x):(x,y)\in E\}$. An $E$-walk from $a_{0}$ to $a_{n}$ is a finite sequence $p=(a_{0}E_{1}a_{1}E_{2}a_{2}\dots E_{n}a_{n})$, where $E_{i}\in\{E,E^{-1}\}$ and $(a_{i-1},a_{i})\in E_{i}$ for all $i$. We say that $p$ is closed if $a_{0}=a_{n}$. The algebraic length of $p$ is the number of occurrences of $E$ minus the number of occurrences of $E^{-1}$. A digraph $\mathbb{G}=(G;E)$ is said to have algebraic length $1$ if there exists a closed $E$-walk of algebraic length $1$. A digraph $\mathbb{G}$ is weakly connected if for all distinct $a,b\in G$ there exists an $E$-walk from $a$ to $b$. A subset $W\subseteq G$ is called a weakly connected component of $\mathbb{G}$ if it is maximal with respect to inclusion such that the digraph $\mathbb{G}|_{W}\coloneqq(W;E|_{W\times W})$ is weakly connected. A fence in $E$ from $a_{0}$ to $a_{2n}$ is an $E$-walk of the form $(a_{0}\dots a_{2n})$ where $(a_{2i-2},a_{2i-1})\in E$ and $(a_{2i-1},a_{2i})\in E^{-1}$ for all $i\leq n$. The vertices $a_{0},a_{2},\dots,a_{2n}$ are called the lower tips of $p$. For $m\in\mathbb{N}$, a fence in $E^{m}$ is also called an $m$-fence in $E$. We say that $\mathbb{G}$ is linked if there exists $m\in\mathbb{N}$ such that for any $a,b\in\textnormal{pr}_{1}(E)$ there exists an $m$-fence in $E$ from $a$ to $b$. A finite smooth digraph is linked if and only if it is weakly connected and has algebraic length $1$ [BK12, Claim 3.8].
A hypergraph is any structure $\mathbb{T}=(T;R)$ containing only one relation $R$, where $R$ is not necessarily binary. Let $n$ be the arity of $R$, and let $\Sigma$ be a subgroup of the symmetric group $\textnormal{Sym}(n)$ on $n$ elements. We say that $\mathbb{T}$ is $\Sigma$-invariant if for every $a\in R$ and $\pi\in\Sigma$ also $(a_{\pi(1)},\dots,a_{\pi(n)})\in R$. A hypergraph is symmetric if it is $\textnormal{Sym}(n)$-invariant, cyclic if it is invariant with respect to the group of cyclic shifts, and $2$-transitive if it is $\Sigma$-invariant for some $2$-transitive group $\Sigma$. A loop is a constant tuple $(a,\dots,a)\in R$.
@@ 2.3. Algebraic
A polymorphism of arity $n\geq 1$ of a structure $\mathbb{A}=(A;\mathcal{R})$ is a mapping $f:A^{n}\to A$ that preserves every $R\in\mathcal{R}$, i.e.  for every $R\in\mathcal{R}$ and any $r_{1},\dots,r_{n}\in R$, the tuple $f(r_{1},\dots,r_{n})$ is contained in $R$. For an $\omega$-categorical structure $\mathbb{A}$, a relation $R$ is pp-definable in $\mathbb{A}$ if and only if it is preserved by all polymorphisms of $\mathbb{A}$ [BN06]. The polymorphism clone of $\mathbb{A}$ is the set of all polymorphism of $\mathbb{A}$, and it is denoted by $\operatorname{\textnormal{Pol}}(\mathbb{A})$. Observe that it indeed forms a clone in the sense of universal algebra, as it is closed under composition and contains all projections.
@@ 2.4. Temporal constraint languages
A temporal relation is a relation first-order definable in $(\mathbb{Q};<)$. A temporal constraint language is a relational structure all of whose relations are temporal. Clearly, the natural action of $\operatorname{\textnormal{Aut}}(\mathbb{Q})\coloneqq\operatorname{\textnormal{Aut}}((\mathbb{Q};<))$ on $\mathbb{Q}^{k}$ has finitely many orbits for every $k\in\mathbb{N}$. The automorphisms of a structure preserve all relations first-order definable in it. As a consequence, the automorphism group of a first-order reduct always contains the original automorphisms. It follows that the first-order reducts of $\omega$-categorical structures, and in particular all temporal constraint languages, are themselves $\omega$-categorical.
Let $F$ be a set of operations on $\mathbb{Q}$. We say that the clone generated by $F$ is the smallest clone of operations $\mathcal{C}$ that contains $F\cup\operatorname{\textnormal{Aut}}(\mathbb{Q})$ and is closed under interpolation; that is, an operation $g$ belongs to $\mathcal{C}$ if and only if for every finite subset $A\subseteq\mathbb{Q}$, there is $f\in\mathcal{C}$ agreeing with $g$ on $A$. In the case that $F$ contains only one operation $f$, we say that $f$ generates $g$. The clones generated by an operation $f$ and a set $F$ of operations are denoted by $\langle f\rangle$ and $\langle F\rangle$, respectively. For $S\subseteq\mathbb{Q}^{k}$ and an operation $f$ on $\mathbb{Q}$, we denote by $\langle S\rangle_{f}$ the smallest subset of $\mathbb{Q}^{k}$ that contains $S$ and is preserved by every operation of $\langle f\rangle.$
@@
As explained in [BK10], such endomorphisms can be constructed inductively. In this construction, one can easily enforce both that any of the endomorphisms has finitely many specified fixed points, or none at all. Note that our definition of $\operatorname{mi}$ differs slightly from the original one in [BK10]. The two operations generate the same clone, and in particular each other. Our definition of $\operatorname{mi}$ is in line with [BR22].
@@
Let $\mathbb{A}$ be a temporal constraint language. If $\mathbb{A}$ is not omni-expressive, then it is preserved by one of $\min$, $\operatorname{mi}$, $\operatorname{mx}$, $\ell\ell$, their duals, or a constant operation, and $\operatorname{\textnormal{CSP}}(\mathbb{A})$ is solvable in polynomial time.
@@
For the proofs of Theorems 1.3 and 1.6, we will additionally draw on two binary operations $\operatorname{\textnormal{lex}}$ and $pp$. As $\ell\ell$ and $\operatorname{dual}(\ell\ell)$ generate $\operatorname{\textnormal{lex}}$, and each of $\min,\operatorname{mi},$ and $\operatorname{mx}$ generates $pp$, it follows that one of $\operatorname{\textnormal{lex}}$, $p$,$p$, and $\operatorname{dual}(pp)$ is contained in the polymorphism clone of every not omni-expressive constraint language. It is worth noting, however, that the presence of $\operatorname{\textnormal{lex}}$ or $pp$ in a polymorphism clone does not, on its own, imply polynomial-time solvability of the underlying template. For example, the temporal constraint language modelling the classical Betweenness problem is preserved by $\operatorname{\textnormal{lex}}$, yet is NP-complete [Opa79].
@@
Let $\operatorname{\textnormal{lex}}:\mathbb{Q}^{2}\to\mathbb{Q}$ be a binary operation satisfying $\operatorname{\textnormal{lex}}(x,y)\leq\operatorname{\textnormal{lex}}(x^{\prime},y^{\prime})$ if and only if
@@
Clearly, such an operation exists. It is easy to see that all operations with these properties generate the same clone [BK10, Observation 1]. As $\ell\ell$ and $\operatorname{dual}(\ell\ell)$ also satisfy them when restricted to $\{(x,y)\in\mathbb{Q}\times\mathbb{Q}\colon x,y<0\}$ and $\{(x,y)\in\mathbb{Q}\times\mathbb{Q}\colon x,y>0\}$, respectively, both of them generate $\operatorname{\textnormal{lex}}$. Note that lex is, by definition, injective.
@@
For every $k\in\mathbb{N}$, we denote by $\sim_{k}$ the orbit-equivalence of $\operatorname{\textnormal{Aut}}(\mathbb{Q})\curvearrowright\mathbb{Q}^{k}$. In other words, for $x,y\in\mathbb{Q}^{k}$ we have $x\sim_{k}y$ if and only if $x_{i}\leq x_{j}\leftrightarrow y_{i}\leq y_{j}$ for all $i,j\leq k$. A pseudo-loop in an $n$-ary relation $R$ on $\mathbb{Q}^{k}$ is a tuple $(a_{1},\dots,a_{n})$ with $a_{1}\sim_{k}\dots\sim_{k}a_{n}.$ The $\sim_{k}$-factor of $R$ is the relation defined on the set of all $\sim_{k}$}-classes of $\mathbb{Q}^{k}$ that contains a tuple $(A_{1},\dots,A_{n})$ of $\sim_{k}$-classes if and only if there exists a tuple $(a_{1},\dots,a_{n})\in R$ such that $a_{i}\in A_{i}$ for every $i\in[n].$ We say that a binary relation $E$ on $\mathbb{Q}^{k}$ has pseudo-algebraic length $1$ if the $\sim_{k}$-factor of $E$ admits a closed walk of algebraic length $1$. Whenever convenient, we naturally consider $n$-tuples on $\mathbb{Q}^{k}$ as tuples on $\mathbb{Q}^{nk}.$
Finally, we note that if $\mathbb{G}$ is pp-interpretable in a temporal constraint language $\mathbb{A}$, then every polymorphism of $\mathbb{A}$ induces a polymorphism of $\mathbb{G}$ [BJK05]: this follows from the fact that we can regard $\mathbb{G}$ as a structure whose domain is a pp-definable subset of some power of $\mathbb{Q}$ factored by a pp-definable equivalence relation, with all its relations being pp-definable as well. In particular, this is true for the automorphisms of $\mathbb{Q}$, hence $\operatorname{\textnormal{Aut}}(\mathbb{Q})\curvearrowright\mathbb{G}$. If $\mathbb{G}=(G;E)$ is a digraph, a pseudo-loop in $\mathbb{G}$ modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$ is an edge whose vertices are contained in the same orbit of $\operatorname{\textnormal{Aut}}(\mathbb{Q})\curvearrowright\mathbb{G}$. We say that $\mathbb{G}$ has pseudo-algebraic length $1$ modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$ if there exists an $E$-walk of algebraic length $1$ so that its start point and its end point are contained in the same $\operatorname{\textnormal{Aut}}(\mathbb{Q})$-orbit. Similarly, a pseudo-loop modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$ in an $n$-ary hypergraph $\mathbb{T}=(T;R)$ that is pp-interpretable in a temporal constraint language is an edge $(a_{1},\dots,a_{n})\in R$ such that all $a_{i}$ belong to the same $\operatorname{\textnormal{Aut}}(\mathbb{Q})$-orbit.
@@ 3. Chasing Orbits
For the proofs of Theorems 1.3 and 1.6, a pseudo-loop is constructed via a recursive procedure that, at each step, aligns the minimal elements of appropriately chosen tuples. We only consider relations on some power of $\mathbb{Q}$ that are preserved by one of $\min,\operatorname{mi},\operatorname{mx},$ or $\ell\ell$. The general case follows easily as discussed in the proof of Theorem 1.3.
@@
Let $k\geq 1$.
  • For every $a\in\mathbb{Q}^{k}$, we denote by $\min(a)$ the minimal entry of $a$. By $\operatorname{\textnormal{minx}}(a)$ we denote the set of all indices $i\leq k$ for which $a_{i}=\min(a)$.
  • For all $a,b\in\mathbb{Q}^{k}$ and $I\subseteq[k]$, we write $a\sim_{I}b$ if $\textnormal{pr}_{I}(a)\sim_{|I|}\textnormal{pr}_{I}(b)$.
@@
    $\textnormal{pr}_{I_{m}(a^{m+1}_{i})}(a^{m+1}_{i})\sim_{m}a^{m}_{i}$, i.e. the $\sim_{m}$-class of the subtuple of $a^{m+1}_{i}$ containing the $m$ smallest entries of $a^{m+1}_{i}$ coincides with the $\sim_{m}$-class of the subtuple of $a^{m}_{i}$ containing all of its $m$ smallest entries.
After at most $k$-many steps, this recursion yields a tuple all of whose coordinates lie in the same orbit of $\operatorname{\textnormal{Aut}}(\mathbb{Q})\curvearrowright\mathbb{Q}^{k}$, i.e. a pseudo-loop in $R$.
@@
Let $a$ be an $n$-tuple on $\mathbb{Q}^{k}$. By $M(a)$ we denote the set consisting of all $i\in[n]$ for which $\min(a_{i})=\min(a)$. We say that $a$ is min-clean if for all $i,j\in M(a)$ we have $\operatorname{\textnormal{minx}}(a_{i})=\operatorname{\textnormal{minx}}(a_{j})$.
The technical part of the proofs of Theorems 1.3 and 1.6 concerns the existence of min-clean tuples, and is deferred to Section 5. In the following, we construct a pseudo-loop under the assumption that min-clean tuples exist. We note an asymmetry between the cases distinguished by Theorem 2.3. When the template is preserved by $\ell\ell$, the existence of min-clean tuples only relies on the polymorphism $\operatorname{\textnormal{lex}}$ (which is generated by $\ell\ell$), while the subsequent construction uses the full strength of $\ell\ell$. In the other cases, the construction proceeds using $pp$ (which alone does not prevent omni-expressivity), and the existence of min-clean tuples is shown by employing the polymorphisms $\min,\operatorname{mi}$, and $\operatorname{mx}$, respectively. The construction is made up of a nested composition of $\ell\ell$ and $pp$, building on the following easy observation: let $a\in\mathbb{Q}^{k}$, $q\in\mathbb{Q}$, and define $J\coloneqq\{i\in[k]:a_{i}\leq q\}$. For every $b\in\mathbb{Q}^{k}$, the $\sim_{k}$-class of the $k$-tuple $pp_{q}(a,b)$ is determined only by the $\sim_{|J|}$-class of the $|J|$-tuple $\textnormal{pr}_{J}(a)$ and the $\sim_{(k-|J|)}$-class of the (k|J|)(k-|J|)-tuple $\textnormal{pr}_{[k]\setminus J}(b)$. The same holds true for the tuple ${\ell\ell}_{q}(a,b)$, provided that $\ker(a)=\ker(b)$.
For a binary operation $f$ and $m\geq 2$, we denote by $f^{[m]}$} the $m$-ary operation
${f^{[m]}}(x_{1},\dots,x_{m})\coloneqq f(x_{1},f(x_{2},\dots,f(x_{m-1},x_{m})\dots)).$
Note that $f^{[m]}$} is lies within every clone containing $f$.

### v2

@@ 2.2. Graph-theoretic
A digraph is a relational structure of the form $\mathbb{G}=(G;E)$, where $E$ is binary. $\mathbb{G}$ is called smooth if $\operatorname{pr}_{1}(E)=\operatorname{pr}_{2}(E)$. In this case, we also write $\operatorname{supp}(\mathbb{G})$ for the set $\operatorname{pr}_{1}(E)$. If $E$ is symmetric, $\mathbb{G}$ is also called a graph. For $m\in\mathbb{N}$, we denote by $E^{m}$ the binary relation containing all tuples $(a_{0},a_{m})\in G\times G$ for which there exist $a_{1},\dots,a_{m-1}$ such that $(a_{i-1},a_{i})\in E$ for all $i\leq m$. We write $E^{-1}$ for the relation $\{(y,x):(x,y)\in E\}$. An $E$-walk from $a_{0}$ to $a_{n}$ is a finite sequence $p=(a_{0}E_{1}a_{1}E_{2}a_{2}\dots E_{n}a_{n})$, where $E_{i}\in\{E,E^{-1}\}$ and $(a_{i-1},a_{i})\in E_{i}$ for all $i$. We say that $p$ is closed if $a_{0}=a_{n}$. The algebraic length of $p$ is the number of occurrences of $E$ minus the number of occurrences of $E^{-1}$. A digraph $\mathbb{G}=(G;E)$ is said to have algebraic length $1$ if there exists a closed $E$-walk of algebraic length $1$. A digraph $\mathbb{G}$ is weakly connected if for all distinct $a,b\in G$ there exists an $E$-walk from $a$ to $b$. A fence in $E$ from $a_{0}$ to $a_{2n}$ is an $E$-walk of the form $(a_{0}\dots a_{2n})$ where $(a_{2i-2},a_{2i-1})\in E$ and $(a_{2i-1},a_{2i})\in E^{-1}$ for all $i\leq n$. The vertices $a_{0},a_{2},\dots,a_{2n}$ are called the lower tips of $p$. For $m\in\mathbb{N}$, a fence in $E^{m}$ is also called an $m$-fence in $E$. We say that $\mathbb{G}$ is linked if there exists $m\in\mathbb{N}$ such that for any $a,b\in\operatorname{pr}_{1}(E)$ there exists an $m$-fence in $E$ from $a$ to $b$. A finite smooth digraph is linked if and only if it is weakly connected and has algebraic length $1$ [BK12, Claim 3.8].
A hypergraph is any structure $\mathbb{T}=(T;R)$ containing only one relation $R$, where $R$ is not necessarily binary. We call elements of $R$ hyperedges. Let $n$ be the arity of $R$, and let $\Sigma$ be a subgroup of the symmetric group $\textnormal{Sym}(n)$ on $n$ elements. We say that $\mathbb{T}$ is $\Sigma$-invariant if for every $a\in R$ and $\pi\in\Sigma$ also $(a_{\pi(1)},\dots,a_{\pi(n)})\in R$. A hypergraph is cyclic if it is invariant with respect to the group of cyclic shifts, and $2$-transitive if it is $\Sigma$-invariant for some $2$-transitive group $\Sigma$. A loop is a constant tuple $(a,\dots,a)\in R$.
@@ 2.3. Algebraic
A polymorphism of arity $n\geq 2$ of a structure $\mathbb{A}=(A;\mathcal{R})$ is a mapping $f:A^{n}\to A$ that preserves every $R\in\mathcal{R}$, i.e. for every $R\in\mathcal{R}$ and any $r_{1},\dots,r_{n}\in R$, the tuple $f(r_{1},\dots,r_{n})$ is contained in $R$. For an $\omega$-categorical structure $\mathbb{A}$, a relation $R$ is pp-definable in $\mathbb{A}$ if and only if it is preserved by all polymorphisms of $\mathbb{A}$ [BN06]. The polymorphism clone of $\mathbb{A}$ is the set of all polymorphism of $\mathbb{A}$, and it is denoted by $\operatorname{Pol}(\mathbb{A})$. Observe that it indeed forms a clone in the sense of universal algebra, as it is closed under composition and contains all projections.
@@ 2.4. Temporal constraint languages
A temporal relation is a relation first-order definable in $(\mathbb{Q};<)$. A temporal constraint language is a relational structure with domain $\mathbb{Q}$ all of whose relations are temporal. Clearly, the natural action of $\operatorname{Aut}(\mathbb{Q})\coloneqq\operatorname{Aut}((\mathbb{Q};<))$ on $\mathbb{Q}^{k}$ has finitely many orbits for every $k\in\mathbb{N}$. The automorphisms of a structure preserve all relations first-order definable in it. As a consequence, the automorphism group of a first-order reduct always contains the original automorphisms. It follows that the first-order reducts of $\omega$-categorical structures, and in particular all temporal constraint languages, are themselves $\omega$-categorical.
Let $F$ be a set of operations on $\mathbb{Q}$. We say that the clone generated by $F$ is the smallest clone of operations $\mathcal{C}$ that contains $F\cup\operatorname{Aut}(\mathbb{Q})$ and is closed under interpolation; that is, an operation $g$ belongs to $\mathcal{C}$ if and only if for every finite subset $A\subseteq\mathbb{Q}$, there is $f\in\mathcal{C}$ agreeing with $g$ on $A$. In the case that $F$ contains only one operation $f$, we say that $f$ generates $g$. The clones generated by an operation $f$ and a set $F$ of operations are denoted by $\langle f\rangle$ and $\langle F\rangle$, respectively. For $S\subseteq\mathbb{Q}^{k}$ and an operation $f$ on $\mathbb{Q}$, we denote by $\langle{S}\rangle_{f}$ the smallest subset of $\mathbb{Q}^{k}$ that contains $S$ and is preserved by every operation of $\langle f\rangle.$
@@
As explained in [BK10], such endomorphisms can be constructed inductively. In this construction, one can easily enforce both that any of the endomorphisms has finitely many specified fixed points, or none at all. All operations that have the listed properties of $\operatorname{mi}$ generate the same clone, and the same holds for $\operatorname{mx}$. Note that our definition of $\operatorname{mi}$ differs slightly from the original one in [BK10]. Operations $\operatorname{mi}$ defined either way generate the same clone, and in particular each other. Our definition of $\operatorname{mi}$ is in line with [BR22].
@@
Let $\mathbb{A}$ be a temporal constraint language. If $\mathbb{A}$ is not omni-expressive, then it is preserved by one of $\min$, $\operatorname{mi}$, $\operatorname{mx}$, $\ell\ell$, their duals, or a constant operation, and $\operatorname{CSP}(\mathbb{A})$ is solvable in polynomial time.
@@
For the proofs of Theorems 1.3 and 1.5, we will additionally draw on two binary operations $\operatorname{lex}$ and $pp$. As $\ell\ell$ and $\operatorname{dual}(\ell\ell)$ generate $\operatorname{lex}$, and each of $\min,\operatorname{mi},$ and $\operatorname{mx}$ generates $pp$, it follows that one of $\operatorname{lex}$, $p$,$p$, and $\operatorname{dual}(pp)$ is contained in the polymorphism clone of every not omni-expressive constraint language. It is worth noting, however, that the presence of $\operatorname{lex}$ or $pp$ in a polymorphism clone does not, on its own, imply polynomial-time solvability of the underlying template. For example, the temporal constraint language modelling the classical Betweenness problem is preserved by $\operatorname{lex}$, yet is NP-complete [Opa79].
@@
Let $\operatorname{lex}:\mathbb{Q}^{2}\to\mathbb{Q}$ be a binary operation satisfying $\operatorname{lex}(x,y)\leq\operatorname{lex}(x^{\prime},y^{\prime})$ if and only if
@@
Clearly, such an operation exists. It is easy to see that all operations with these properties generate the same clone [BK10, Observation 1]. As $\ell\ell$ and $\operatorname{dual}(\ell\ell)$ also satisfy them when restricted to $\{(x,y)\in\mathbb{Q}\times\mathbb{Q}\colon x,y<0\}$ and $\{(x,y)\in\mathbb{Q}\times\mathbb{Q}\colon x,y>0\}$, respectively, both of them generate $\operatorname{lex}$. Note that lex is, by definition, injective.
@@
For every $k\in\mathbb{N}$, we denote by $\sim_{k}$ the orbit-equivalence of $\operatorname{Aut}(\mathbb{Q})\curvearrowright\mathbb{Q}^{k}$. In other words, for $x,y\in\mathbb{Q}^{k}$ we have $x\sim_{k}y$ if and only if $x_{i}\leq x_{j}\leftrightarrow y_{i}\leq y_{j}$ for all $i,j\leq k$. A pseudo-loop in an $n$-ary relation $R$ on $\mathbb{Q}^{k}$ is a tuple $(a_{1},\dots,a_{n})$ with $a_{1}\sim_{k}\dots\sim_{k}a_{n}.$ The $\sim_{k}$-factor of $R$ is the relation defined on the set of all $\sim_{k}$}-classes of $\mathbb{Q}^{k}$ that contains a tuple $(A_{1},\dots,A_{n})$ of $\sim_{k}$-classes if and only if there exists a tuple $(a_{1},\dots,a_{n})\in R$ such that $a_{i}\in A_{i}$ for every $i\in[n].$ We say that a binary relation $E$ on $\mathbb{Q}^{k}$ has pseudo-algebraic length $1$ if the $\sim_{k}$-factor of $E$ admits a closed walk of algebraic length $1$. Whenever convenient, we naturally consider $n$-tuples on $\mathbb{Q}^{k}$ as tuples on $\mathbb{Q}^{nk}.$
Finally, we note that if $\mathbb{G}$ is pp-interpretable in a temporal constraint language $\mathbb{A}$, then every polymorphism of $\mathbb{A}$ induces a polymorphism of $\mathbb{G}$ [BJK05]: this follows from the fact that we can regard $\mathbb{G}$ as a structure whose domain is a pp-definable subset of some power of $\mathbb{Q}$ factored by a pp-definable equivalence relation, with all its relations being pp-definable as well. In particular, this is true for the automorphisms of $\mathbb{Q}$, hence $\operatorname{Aut}(\mathbb{Q})\curvearrowright\mathbb{G}$. If $\mathbb{G}=(G;E)$ is a digraph, a pseudo-loop in $\mathbb{G}$ modulo $\operatorname{Aut}(\mathbb{Q})$ is an edge whose vertices are contained in the same orbit of $\operatorname{Aut}(\mathbb{Q})\curvearrowright\mathbb{G}$. We say that $\mathbb{G}$ has pseudo-algebraic length $1$ modulo $\operatorname{Aut}(\mathbb{Q})$ if there exists an $E$-walk of algebraic length $1$ so that its start point and its end point are contained in the same $\operatorname{Aut}(\mathbb{Q})$-orbit. Similarly, a pseudo-loop modulo $\operatorname{Aut}(\mathbb{Q})$ in an $n$-ary hypergraph $\mathbb{T}=(T;R)$ that is pp-interpretable in a temporal constraint language is an edge $(a_{1},\dots,a_{n})\in R$ such that all $a_{i}$ belong to the same $\operatorname{Aut}(\mathbb{Q})$-orbit.
@@ 3. Chasing Orbits
For the proofs of Theorems 1.3 and 1.5, a pseudo-loop is constructed via a recursive procedure that, at each step, aligns the minimal elements of appropriately chosen tuples. We only consider relations on some power of $\mathbb{Q}$ that are preserved by one of the operations $\min,\operatorname{mi},\operatorname{mx},$ or $\ell\ell$. The general case for relations pp-interpretable in temporal constraint languages easily reduces to this setting, as discussed in the proof of Theorem 1.3. In what follows, we fix an integer $k\geq 2$.
@@
x
  • For every $a\in\mathbb{Q}^{k}$, we denote by $\min(a)$ the minimal entry of $a$. By $\operatorname{minx}(a)$ we denote the set of all indices $i\leq k$ for which $a_{i}=\min(a)$.
  • For all $a,b\in\mathbb{Q}^{k}$ and $I\subseteq[k]$, we write $a\sim_{I}b$ if $\operatorname{pr}_{I}(a)\sim_{|I|}\operatorname{pr}_{I}(b)$.
@@
    $\operatorname{pr}_{I_{m}(a^{m+1}_{i})}(a^{m+1}_{i})\sim_{m}\operatorname{pr}_{I_{m}(a^{m}_{i})}(a^{m}_{i})$, i.e. the $\sim_{m}$-class of the subtuple of $a^{m+1}_{i}$ containing the $m$ smallest entries of $a^{m+1}_{i}$ coincides with the $\sim_{m}$-class of the subtuple of $a^{m}_{i}$ containing all of its $m$ smallest entries.
After at most $k$-many steps, this recursion yields a tuple all of whose coordinates lie in the same orbit of $\operatorname{Aut}(\mathbb{Q})\curvearrowright\mathbb{Q}^{k}$, i.e. a pseudo-loop in $R$.
@@
Let $a$ be an $n$-tuple on $\mathbb{Q}^{k}$. By $M(a)$ we denote the set consisting of all $i\in[n]$ for which $\min(a_{i})=\min(a)$. We say that $a$ is min-clean if for all $i,j\in M(a)$ we have $\operatorname{minx}(a_{i})=\operatorname{minx}(a_{j})$.
The technical part of the proofs of Theorems 1.3 and 1.5 concerns the existence of min-clean tuples, and is deferred to Section 6. In the following, we construct a pseudo-loop under the assumption that min-clean tuples exist. We note an asymmetry between the cases distinguished by Theorem 2.3. When the template is preserved by $\ell\ell$, the existence of min-clean tuples only relies on the polymorphism $\operatorname{lex}$ (which is generated by $\ell\ell$), while the subsequent construction uses the full strength of $\ell\ell$. In the other cases, the construction proceeds using $pp$ (which alone does not prevent omni-expressivity), and the existence of min-clean tuples is shown by employing the polymorphisms $\min,\operatorname{mi}$, and $\operatorname{mx}$, respectively. The construction is made up of a nested composition of $\ell\ell$ and $pp$, building on the following easy observation: let $a\in\mathbb{Q}^{k}$, $q\in\mathbb{Q}$, and define $J\coloneqq\{i\in[k]:a_{i}\leq q\}$. For every $b\in\mathbb{Q}^{k}$, the $\sim_{k}$-class of the $k$-tuple $pp_{q}(a,b)$ is determined only by the $\sim_{|J|}$-class of the $|J|$-tuple $\operatorname{pr}_{J}(a)$ and the $\sim_{(k-|J|)}$-class of the (k|J|)(k-|J|)-tuple $\operatorname{pr}_{[k]\setminus J}(b)$. The same holds true for the tuple ${\ell\ell}_{q}(a,b)$, provided that $\ker(a)=\ker(b)$. To simplify notation, for a binary operation $f$ and $m\geq 2$, we denote by $f^{[m]}$} the $m$-ary operation defined as ${f^{[m]}}(x_{1},\dots,x_{m})\coloneqq f(x_{1},f(x_{2},\dots,f(x_{m-1},x_{m})\dots)).$ Note that $f^{[m]}$} is lies within every clone containing $f$.